OpenStudy (anonymous):

Use the fundamental identities to simplify this expression :csc x/cotxsecx PLEASE SHOW ME STEP BY STEP HOW TO GET THE ANSWER!

OpenStudy (solomonzelman):

here are the rules: 1) \(\large\color{black}{ \csc(x)=\frac{\LARGE 1}{\LARGE \sin(x)} }\) 2) \(\large\color{black}{ \sec(x)=\frac{\LARGE 1}{\LARGE \cos(x)} }\) 3) \(\large\color{black}{ \cot(x)=\frac{\LARGE \cos(x)}{\LARGE \sin(x)} }\)

OpenStudy (solomonzelman):

And the problem is, \(\large\color{black}{ \frac{ \LARGE \csc(x)}{\LARGE \cot(x)\sec(x)} }\) correct?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok i got to that step

OpenStudy (anonymous):

Do you just cancel them out now?

OpenStudy (solomonzelman):

Which step did you get? I haven't posted any "steps" I posted the rules and the problem.

OpenStudy (anonymous):

The seperation using quotient identity and reciprocal

OpenStudy (solomonzelman):

This is what I would do, \(\large\color{black}{ \frac{ \LARGE \csc(x)}{\LARGE \cot(x)\sec(x)} }\) applying rule 3, we get, \(\large\color{black}{ \frac{ \LARGE \csc(x)}{\frac{\huge\cos(x)}{\huge \sin(x)}\LARGE \sec(x)} }\) applying rule 2 we get, \(\large\color{black}{ \frac{ \LARGE \csc(x)}{\frac{\huge\cos(x)}{\huge \sin(x)}~\times~\frac{\huge1}{\huge \cos(x)}} }\)

OpenStudy (solomonzelman):

what cancels?

OpenStudy (anonymous):

the cosines

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

\(\large\color{black}{ \frac{ \LARGE \csc(x)}{\frac{\huge1}{\huge \sin(x)}} }\)

OpenStudy (solomonzelman):

and this is equal to?

OpenStudy (solomonzelman):

(Rule 1)

OpenStudy (anonymous):

1

OpenStudy (solomonzelman):

Yup !

OpenStudy (anonymous):

1/sin*1/sin

OpenStudy (anonymous):

thank you

OpenStudy (solomonzelman):

the answer is 1

OpenStudy (anonymous):

Could you help me with one more

OpenStudy (solomonzelman):

yw

OpenStudy (solomonzelman):

I can try, lol;)

OpenStudy (anonymous):

\[\csc^2\theta-\cot^2\theta/\sec\]

OpenStudy (solomonzelman):

just use the same rules, to re-write everything in terms of sine and cosine

OpenStudy (solomonzelman):

Wait, the numerator is csc^2x-cot^2x ?

OpenStudy (anonymous):

yea

OpenStudy (solomonzelman):

csc^2x = cot^2x+1 for all x values.

OpenStudy (solomonzelman):

So, re-write the numerator, what do you get?

OpenStudy (anonymous):

cot^2x+1-cot^2x

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

and that is equal to?

OpenStudy (anonymous):

1/secx

OpenStudy (anonymous):

i mean 1 and then put it over the denominator

OpenStudy (solomonzelman):

yes, and 1/sec(x) is?

OpenStudy (anonymous):

cos x thank you soooo much!

OpenStudy (solomonzelman):

yes, and 1/sec(x) is?

OpenStudy (solomonzelman):

yes yes, sorry, cos(x)....

OpenStudy (solomonzelman):

Didn't see your freply b4...

OpenStudy (anonymous):

ur fine lol